# Quadrature Coherence Scale of Linear Combinations of Gaussian Functions in Phase Space

### v1

@@ 6.2 GKP states
where $|\cdot\rangle_{q}$ denotes an eigenstate of the position quadrature. Due to this gridlike property, GKP states can be used to correct displacement errors, making them essential resources for continous-variable quantum computation [85, 86, 87]
The ideal GKP state has infinite energy and cannot be normalizable. One thus needs to define finite-energy GKP states in order to deal with them. One option is to apply a Fock damping operator $E(\epsilon)={\mathrm{e}}^{-\epsilon\hat{n}}$ with $\epsilon>0$ [45, 88]. It can then be shown that the Wigner function of the GKP states can be written as a linear combination of Gaussian functions in phase space, as defined in Eq. (11). Let us introduce
@@
$\boldsymbol{\mu}_{m}=\frac{\sqrt{\pi}}{2}\begin{pmatrix}\text{sech}(\epsilon)(2k+2l+s+t)\\ i\,\text{csch}(\epsilon)(-2k+2l+s-t),\end{pmatrix},\ \text{and}$ (30)
@@
Figure 1: Plot of the QCS (squared) of a GKP state as a function of the damping operator $\epsilon$. The green dashed curve shows the general behaviour of the QCS.
6.3 Evolution of the QCS through a loss channel
It is essential to consider loss in any actual experiment. In the domain of state generation, if the goal is to generate a nonclassical state, it is important to understand how the nonclassicality, e.g., the QCS, evolves with losses. A simple model of loss is the loss channel whose action in the phase space is described by Eq. (5). As mentioned in Sec. 4, since a loss channel is described by a Gaussian transformation, when the Wigner function of the input state can be written as a sum of Gaussian functions, so will the Wigner function of the output state. The coefficient $c_{m}$ in Eqs. (11) remains the same and the covariance matrices and mean values are updated as follows:
@@ 6.3 Evolution of the QCS through a loss channel
Loss is the dominant source of errors in photon-based protocols and can arise from multiple physical mechanisms including imperfect coupling and detector inefficiencies. Fortunately, loss is additive, such that the effects of disparate loss processes can be aggregated into a single parameter $\eta$. This provides a microscopic description that, as expected, reduces the average energy in a beam by a multiplicative factor of $\eta$. Without loss, one might expect fault-tolerant quantum computation with continuous variables to already be prevalent.
@@
6.3.1 Cat states in a loss channel
@@
As already shown in [37], we observe that the decoherence rate grows quadratically with the QCS (and quadratically with the energy) of the cat state: the larger the nonclassicality, the quicker we lose it.
It is interesting to note that, no matter the value of $\alpha$, the QCS of the states reaches 1 (which is the nonclassicality threshold) for $\eta=1/2$. This phenomenon was already observed in Ref. [40].
@@
6.3.2 GKP states in a loss channel
@@
6.3.3 Squeezed state in a loss channel
As another example, it is also possible to compute the QCS of a lossy squeezed vacuum state $S(r)|0\rangle$, where $S(r)={\mathrm{e}}^{\frac{1}{2}(\bar{r}a^{2}-ra^{{\dagger}2})}$ is the squeezing operator. We obtain (this result was already derived in Ref. [40]):
@@ 6.3.3 Squeezed state in a loss channel
Figure 4: Comparison of the loss (parameterized by $\eta$) of nonclassicality, as measured by the QCS for a cat state (plain line), a GKP state (dashed line), and a squeezed state (dotted line). The orange curves represent states starting with ${\mathcal{C}}^{2}\approx 16.5$ (i.e. $\alpha\approx 2.8$, $\epsilon\approx 0.05$ and $r=1.7$) and blue curves states starting with ${\mathcal{C}}^{2}\approx 4.4$ (i.e. $\alpha\approx 1.3$, $\epsilon\approx 0.2$ and $r=1.1$).
@@
6.4 Breeding
Several schemes exist that strive to efficiently generate GKP states in optical setups [90, 91, 92, 86, 88, 93, 94, 95, 96, 97, 98, 99]. The breeding protocol [100, 14, 101, 50, 15] is a procedure that gradually generates (or “breeds”) GKP states by impinging squeezed cat states $\left(1+D(\alpha)\right)S(r)|0\rangle$ on a 50:50 beam splitter and measuring the $p$-quadrature of one output mode. By post-selecting on specific values of the $p$-quadrature, one can obtain a GKP state in the other output mode of the beam splitter. This procedure can be repeated several times in order to increase the fidelity of the output state. In one recent protocol, the vast majority of $p$-measurement values suffice for generating GKP states, singling out GKP state as “nonclassical fixed points” of linear optics. The result is an approximation of a GKP state with $\sqrt{2}\alpha$ spacing and ${\mathrm{e}}^{r}$ squeezing.
In what is called a slow breeding protocol, the output of the first round of breeding is inserted again with another squeezed cat state of displacement $\alpha/\sqrt{2}$ at the second input, and the $p$-quadrature of one of the outputs is once again measured. This continues, where in the $M$th round the input squeezed cat state is of the form $\left(1+D(\alpha/\sqrt{2^{M-1}})\right)S(r)|0\rangle$. More details can be found in Ref. [15]. Note that, at each round, one had to do a post-selection on the value of the $p$-quadrature. As in Refs. [15, 50, 100], we choose here to post-select on $p=0$. Nevertheless, keep in mind that the variance of the probability distribution of measuring $p=0$ scales as ${\mathrm{e}}^{2r}$.
@@ 6.4 Breeding
One important question is to understand how many rounds of breeding one needs to get a GKP state with high enough fidelity. And how can this fidelity be measured? We suggest here to look at the QCS. Indeed, we know that the QCS of a cat state grows with its amplitude $\alpha$ and we thus expect the QCS to increase through the breeding protocol. Since the output can be written as a sum of Gaussian functions and the breeding circuit is a Gaussian transformation, the QCS of the output can be easily computed with the equations presented in Sec. 5.
@@
The output state will then have the form [15]
@@
Figure 6: Wigner functions of a GKP state generated after M=0,1,2,3,4 rounds of the slow breeding protocol with final spacing of $2\sqrt{\pi}$ and squeezing $r=-\ln(0.2)$. Note that at each step, we applied a final displacement of $(M+1)\sqrt{\pi/2}$ in the $x$-direction to center the state on 00.
@@
It has the Wigner function of a normalized Gaussian of covariance matrix $\boldsymbol{\gamma}$ and mean value $\boldsymbol{\mu}$ given by (see Appendix A of Ref. [45])
@@
Using the equations of Sec. 5, we can compute the QCS at each round. Note that $M=0$ corresponds to the input squeezed cat state and the QCS is given by
@@
In Fig. 7, we compare the QCS of a GKP state as defined in Sec. 6.2 with the QCS of the output state of $M$ rounds of the (slow) breeding protocol. We compare two cases: $r=-\ln(0.2)$ (blue line) as in Ref. [15] and $r=-\ln(0.3)$ (orange line). We observe that the QCS increases rapidly and reaches the ”target” value (i.e. the GKP state as described in Sec. 6.2 with $\tanh\epsilon={\mathrm{e}}^{-2r}$) after only a few rounds. But more importantly, we see that the QCS increases with each round of breeding and thus tends to infinity, as we would expect from a perfect GKP state. This is in contrast with the GKP state as described by the damping operator, but can be explained by looking at the Wigner function.
@@
In Fig. 8, we plotted a slice of the Wigner function $W(x,p=0)$ of a GKP state ($\epsilon\approx 0.09$) and the output state of 5 rounds of the breeding protocol ($r=-\ln(0.3)\approx 1.2$). As we can see, the spacing of each Gaussian peak is the same, but the Gaussian envelope is not. The variance of the Gaussian envelope of the GKP state as described by the damping operator is smaller, so that the total energy of the state as well as the QCS are smaller. On the other hand, after each round of the breeding protocol, the Wigner function has more and more non-zero peaks further from the origin so that both energy and QCS tend to infinity when the number of rounds $M$ grows. This explains why the QCS of the breeding protocol does not tend to the value of the QCS of the targeted GKP state as described by the damping operator in Sec. 6.2.
Figure 8: Comparison of the Wigner function $W(x,p=0)$ of a GKP state (dashed blue line and the output state of 5 rounds of the breeding protocol (plain orange line).
@@
7 Conclusion
@@
This paper presented a general formulation of the QCS of mixed quantum states whose Wigner functions can be expressed as a linear combination of Gaussian functions. This family of states encompasses important non-Gaussian states like Schrödinger cat states and GKP states, but also the output of a breeding protocol or, more generally, all resulting states of Gaussian operations applied to any states of such a form. Note that, in this work, we only considered transformations that are deterministic Gaussian completely positive and trace-preserving maps, but the class of Gaussian operations falling into this formalism is much larger and one could consider conditional dynamics (when the measurement of some modes updates the remaining modes), which shows the versatility and applicability of the proposed framework. With such a transformation, the number as well as the weight of each coefficient $c_{m}$ in Eq. (11) would change. More details can be found in Ref. [45].
@@
Loss as described by a loss channel falls into this framework, which allowed us to assess the scalability of nonclassicality, as measured by the QCS, with loss. As expected, for all states, the QCS decreases when loss increases, as long as the loss is less than 50%. At this exact point, the QCS of all our examples of pure states reached a value of exactly 1, which leads us to conjecture that this is true for all pure states. We extend the conjecture by suggesting that all states lose their nonclassicality at the latest after 50% loss, a statement which is supported by all Wigner functions becoming strictly non-negative at the 50%-loss mark.
@@ A.2 QCS of a Fock state after 50% loss
We know from Ref. [89] that the Wigner function of the output state of the loss channel is proportional to the Q-function of the input state (see Eq. (36)). The Q-function of a Fock state $|k\rangle$ is given by [102]

### v3

@@ VI.2 GKP states
where $|\cdot\rangle_{q}$ denotes an eigenstate of the position quadrature. Due to this gridlike property, GKP states can be used to correct displacement errors, making them essential resources for continuous-variable quantum computation Grimsmo2021 ; Tzitrin.PRXQuantum.2.040353 ; lachancequirion2023autonomous
The ideal GKP state has infinite energy and cannot be normalizable. One thus needs to define finite-energy GKP states in order to deal with them. One option is to apply a Fock damping operator $E(\epsilon)={\mathrm{e}}^{-\epsilon\hat{n}}$ with $\epsilon>0$ XanaduSumGaussian ; XanaduComputationWithGKPstates . It can then be shown that the Wigner function of the GKP states can be written as a linear combination of Gaussian functions in phase space, as defined in Eq. (11). Let us introduce
@@
$\boldsymbol{\mu}_{m}=\frac{\sqrt{\pi}}{2}\begin{pmatrix}\text{sech}(\epsilon)(2k+2l+s+t)\\ i\,\text{csch}(\epsilon)(-2k+2l+s-t),\end{pmatrix},\ \text{and}$
@@
Figure 1: Plot of the QCS (squared) of a GKP state as a function of the damping operator $\epsilon$. The green dashed curve shows the general behaviour of the QCS. All units here and in all subsequent figures are dimensionless.
VI.3 Evolution of the QCS through a loss channel
It is essential to consider loss in any actual experiment. In the domain of state generation, if the goal is to generate a nonclassical state, it is important to understand how the nonclassicality, e.g., the QCS, evolves with losses. A simple model of loss is the loss channel whose action in the phase space is described by Eq. (5). As mentioned in Sec. IV, since a loss channel is described by a Gaussian transformation, when the Wigner function of the input state can be written as a sum of Gaussian functions, so will the Wigner function of the output state. The coefficient $c_{m}$ in Eqs. (11) remains the same and the covariance matrices and mean values are updated as follows:
@@ VI.3 Evolution of the QCS through a loss channel
Loss is the dominant source of errors in photon-based protocols and can arise from multiple physical mechanisms including imperfect coupling and detector inefficiencies. It naturally arose in the reported measurements of the QCS on a cloud quantum computer Goldbergetal23 and similarly arises in many real-life experiments with quantum states PhysRevA.76.032309 ; PhysRevA.82.021801 ; Sperlingetal2015 ; Ferrettietal2024 . Fortunately, loss is additive, such that the effects of disparate loss processes can be aggregated into a single parameter $\eta$. This provides a microscopic description that, as expected, reduces the average energy in a beam by a multiplicative factor of $\eta$. Without loss, one might expect fault-tolerant quantum computation with continuous variables to already be prevalent.
@@
VI.3.1 Cat states in a loss channel
@@
As already shown in Hertz , we observe that the decoherence rate grows quadratically with the QCS (and quadratically with the energy) of the cat state: the larger the nonclassicality, the quicker we lose it.
It is interesting to note that, no matter the value of $\alpha$, the QCS of the states reaches 1 (which is the nonclassicality threshold) for $\eta=1/2$. This phenomenon was already observed in Ref. Goldbergetal23 .
@@
VI.3.2 GKP states in a loss channel
@@
VI.3.3 Squeezed state in a loss channel
As another example, it is also possible to compute the QCS of a lossy squeezed vacuum state $S(r)|0\rangle$, where $S(r)={\mathrm{e}}^{\frac{1}{2}(\bar{r}a^{2}-ra^{{\dagger}2})}$ is the squeezing operator. We obtain (this result was already derived in Ref. Goldbergetal23 ):
@@ VI.3.3 Squeezed state in a loss channel
Figure 4: Comparison of the loss (parameterized by $\eta$) of nonclassicality, as measured by the QCS for a cat state (plain line), a GKP state (dashed line), and a squeezed state (dotted line). The orange curves (highest three at far right) represent states starting with ${\mathcal{C}}^{2}\approx 16.5$ (i.e. $\alpha\approx 2.8$, $\epsilon\approx 0.05$ and $r=1.7$) and blue curves (lowest three at far right) states starting with ${\mathcal{C}}^{2}\approx 4.4$ (i.e. $\alpha\approx 1.3$, $\epsilon\approx 0.2$ and $r=1.1$).
@@
VI.4 Breeding
Several schemes exist that strive to efficiently generate GKP states in optical setups Menicucci2014 ; Eatonetal2019 ; Suetal2019 ; Tzitrin.PRXQuantum.2.040353 ; XanaduComputationWithGKPstates ; Eatonetal2022 ; Takaseetal2023 ; Yanagimotoetal2023 ; Konnoetal2023arxiv ; Crescimannaetal2023arxiv ; AntenehPfister2023arxiv ; Takaseetal2024arxiv . The breeding protocol Breeding_Etesse ; Breeding_Etesse_2 ; Breeding_Sychev ; Breeding_Vasconcelos ; Breeding is a procedure that gradually generates (or “breeds”) GKP states by impinging squeezed cat states $\left(1+D(\alpha)\right)S(r)|0\rangle$ on a 50:50 beam splitter and measuring the $p$-quadrature of one output mode. By post-selecting on specific values of the $p$-quadrature, one can obtain a GKP state in the other output mode of the beam splitter. This procedure can be repeated several times in order to increase the fidelity of the output state. In one recent protocol, the vast majority of $p$-measurement values suffice for generating GKP states, singling out GKP states as “nonclassical fixed points” of linear optics. The result is an approximation of a GKP state with $\sqrt{2}\alpha$ spacing and ${\mathrm{e}}^{r}$ squeezing.
In what is called a slow breeding protocol, the output of the first round of breeding is inserted again with another squeezed cat state of displacement $\alpha/\sqrt{2}$ at the second input, and the $p$-quadrature of one of the outputs is once again measured. This continues, where in the $M$th round the input squeezed cat state is of the form $\left(1+D(\alpha/\sqrt{2^{M-1}})\right)S(r)|0\rangle$. More details can be found in Ref. Breeding . Note that, at each round, one had to do a post-selection on the value of the $p$-quadrature. As in Refs. Breeding ; Breeding_Vasconcelos ; Breeding_Etesse , we choose here to post-select on $p=0$. Nevertheless, keep in mind that the variance of the probability distribution of measuring $p=0$ scales as ${\mathrm{e}}^{2r}$.
@@ VI.4 Breeding
One important question is to understand how many rounds of breeding one needs to get a GKP state with high enough fidelity. And how can this fidelity be measured? We suggest here to look at the QCS. Indeed, we know that the QCS of a cat state grows with its amplitude $\alpha$ and we thus expect the QCS to increase through the breeding protocol. Since the output can be written as a sum of Gaussian functions and the breeding circuit is a Gaussian transformation, the QCS of the output can be easily computed with the equations presented in Sec. V.
@@
The output state will then have the form Breeding
@@
Figure 6: Wigner functions of a GKP state generated after M=0,1,2,3,4 rounds of the slow breeding protocol with final spacing of $2\sqrt{\pi}$ and squeezing $r=-\ln(0.2)$. Note that at each step, we applied a final displacement of $(M+1)\sqrt{\pi/2}$ in the $x$-direction to center the state on 00. Colour bar indicates the magnitude and sign of the Wigner function.
@@
It has the Wigner function of a normalized Gaussian of covariance matrix $\boldsymbol{\gamma}$ and mean value $\boldsymbol{\mu}$ given by (see Appendix A of Ref. XanaduSumGaussian )
@@
Using the equations of Sec. V, we can compute the QCS at each round. Note that $M=0$ corresponds to the input squeezed cat state and the QCS is given by
@@
In Fig. 7, we compare the QCS of a GKP state as defined in Sec. VI.2 with the QCS of the output state of $M$ rounds of the (slow) breeding protocol. We compare two cases: $r=-\ln(0.2)$ (blue line) as in Ref. Breeding and $r=-\ln(0.3)$ (orange line). We observe that the QCS increases rapidly and reaches the ”target” value (i.e. the GKP state as described in Sec. VI.2 with $\tanh\epsilon={\mathrm{e}}^{-2r}$) after only a few rounds. But more importantly, we see that the QCS increases with each round of breeding and thus tends to infinity, as we would expect from a perfect GKP state. This is in contrast with the GKP state as described by the damping operator, but can be explained by looking at the Wigner function.
@@
In Fig. 8, we plotted a slice of the Wigner function $W(x,p=0)$ of a GKP state ($\epsilon\approx 0.09$) and the output state of 5 rounds of the breeding protocol ($r=-\ln(0.3)\approx 1.2$). As we can see, the spacing of each Gaussian peak is the same, but the Gaussian envelope is not. The variance of the Gaussian envelope of the GKP state as described by the damping operator is smaller, so that the total energy of the state as well as the QCS are smaller. On the other hand, after each round of the breeding protocol, the Wigner function has more and more non-zero peaks further from the origin so that both energy and QCS tend to infinity when the number of rounds $M$ grows. This explains why the QCS of the breeding protocol does not tend to the value of the QCS of the targeted GKP state as described by the damping operator in Sec. VI.2.
Figure 8: Comparison of the Wigner function $W(x,p=0)$ of a GKP state (dashed blue line) and the output state of 5 rounds of the breeding protocol (plain orange line).
@@
VII Other measures for sums of Gaussian functions
Suppose we restrict our states to those whose Wigner functions can be written as sums of Gaussian functions. Does that simplify the computation of other measures of quantumness?
The general answer is negative. Consider computation of the quantum Fisher information for a state’s sensitivity to displacements, averaged over all displacement directions. More quantumness means higher average quantum Fisher information, which is a metric specifically tied to the usefulness of the state for metrology. For pure states, this is simply proportional to the average number of photons in the state, the same as the QCS Goldbergetal2020extremal . For mixed states $\rho$, such as a cat state subject to loss, the quantum Fisher information is given by SidhuKok2020
$\mathsf{Q}_{\rho}=-\int_{0}^{\infty}\mathrm{d}s\mathrm{Tr}\{([\hat{x},\rho]{\mathrm{e}}^{-\rho s})^{2}+([\hat{p},\rho]{\mathrm{e}}^{-\rho s})^{2}\}.$ (40)
If not for the factors ${\mathrm{e}}^{-\rho s}$ and the integral over $s$, this would look very similar to the QCS (hence their agreement for pure states); that $\rho$ has a Wigner function that is a linear combination of Gaussians does not simplify this formula in any known way.
One metric that is easy to compute for states that are linear combinations of Gaussians is the total noise; i.e., the trace of the state’s covariance matrix. Unfortunately, this quantity tends to be most useful for pure states, where it is equal to the QCS, but it becomes less insightful for mixed states, where expressions like the quantum Fisher information $\mathsf{Q}_{\rho}$ are more useful and yet do not inherit the ease of computation with linear combinations of Gaussians.
Next, consider computation of a state’s negative Wigner volume
$\mathsf{N}_{\rho}=\int\mathrm{d}^{2}\mathbf{r}\frac{|W(\mathbf{r})|-W(\mathbf{r})}{2}.$ (41)
This formula is straightforward to manipulate when the state is subject to loss, as the Wigner function’s transformation is known [see Eq. (31)]. However, computing the absolute value $|W(\mathbf{r})|$ for a linear combination of Gaussian functions is challenging because each coefficient in the linear combination is, in general, complex. Moreover, the Gaussian functions themselves may be complex, and adding loss changes those complex functions, even if the complex coefficients $c_{m}$ are unchanged by loss.
As for distance to classical states, it is not known whether there exists a faster optimization procedure for states whose Wigner functions are linear combinations of Gaussians. It is indeed the case that the classical states belong to this category, but that does not imply that the “straightest path” from a given state to a given classical state is through the set of states whose Wigner functions are combinations of Gaussians. New optimizations are required for each state and each set of parameters such as loss, so this special class of states does not ameliorate the prohibitiveness of distance measures for straightforward computations.
With that all in mind, we compare the behaviour of $\mathsf{N}_{\rho}$ to ${\mathcal{C}}^{2}$ for the particular class of lossy cat states in Fig. 9. The surprising features that we learned from the QCS again feature in the Wigner negativity, showcasing a correlation in their ability to find physical properties. Nonetheless, $\mathsf{N}$ is not known to bound the distance to classical states and thus, alone, could not be interpreted in any other way than witnessing nonclassicality. In addition, the computation of the QCS was rapid with our formalism while the computation of $\mathsf{N}_{\rho}$ required a costly integration over phase space for each point on the curve, as would be required for computing the negativity of any filtered quasiprobability distribution Tan2020 . That similar information can be gleaned from each confirms the usefulness of our results.
Figure 9: Plot of $\mathsf{N}$, the negative volume of the Wigner function of a cat state going through a loss channel, as functions of $\alpha$ and $\eta$. The peak in $\mathsf{N}$ versus $\alpha$ shifts with $\eta$ in the same way as does the QCS in Fig. 2.
VIII Conclusion
@@
This paper presented a general formulation of the QCS of mixed quantum states whose Wigner functions can be expressed as a linear combination of Gaussian functions. This family of states encompasses important non-Gaussian states like Schrödinger cat states and GKP states, but also the output of a breeding protocol or, more generally, all resulting states of Gaussian operations applied to any states of such a form. Note that, in this work, we only considered transformations that are deterministic Gaussian completely positive and trace-preserving maps, but the class of Gaussian operations falling into this formalism is much larger and one could consider conditional dynamics (when the measurement of some modes updates the remaining modes), which shows the versatility and applicability of the proposed framework. With such a transformation, the number as well as the weight of each coefficient $c_{m}$ in Eq. (11) would change. More details can be found in Ref. XanaduSumGaussian .
The methods presented here could be applied to states whose Wigner functions are more complicated than linear combinations of Gaussian functions, with some modification. If some of the functions in the linear combination are given by $f(\mathbf{r})G(\mathbf{r})$ for some polynomial $f(\mathbf{r})$, there are closed-form solutions for each Gaussian integral for each monomial in the polynomials. This happens, for example, when a Gaussian state is acted on by a polynomial function of creation and annihilation operators. The expressions become more complicated as the degrees of the polynomials increase, especially when faced with linear combinations of different polynomials, but the crux of the methods here that rely on Gaussian integration can be broadly applied.
@@
Loss as described by a loss channel falls into this framework, which allowed us to assess the scalability of nonclassicality, as measured by the QCS, with loss. As expected, for all states, the QCS decreases when loss increases, as long as the loss is less than 50%. At this exact point, the QCS of all our examples of pure states reached a value of exactly 1, which leads us to conjecture that this is true for all pure states. We extend the conjecture by suggesting that all states lose their nonclassicality (as measured by the QCS) at the latest after 50% loss, a statement which is supported by all Wigner functions becoming strictly non-negative at the 50%-loss mark.
@@ A.2 QCS of a Fock state after 50% loss
We know from Ref. RadimFilip that the Wigner function of the output state of the loss channel is proportional to the Q-function of the input state (see Eq. (31)). The Q-function of a Fock state $|k\rangle$ is given by linowski2023relating